OpenStudy (anonymous):

A local pizzeria offers 11 toppings for their pizzas and you can choose any 5 of them for one fixed price. How many different types of pizzas can you order with 5 toppings? A. 462 B. 120 C. 55,440 D. 332,640

OpenStudy (anonymous):

@ganeshie8

ganeshie8 (ganeshie8):

you got 11 toppings, u wanto pick 5 from them

ganeshie8 (ganeshie8):

# of ways of pickign 5 objects from 11 objects : \(\large ^{11}C_5\)

ganeshie8 (ganeshie8):

knw how to evaluate ?

OpenStudy (anonymous):

I'm not sure but is it like 11*12*13*14*15

ganeshie8 (ganeshie8):

something like that... but u need to divide 5! also

OpenStudy (anonymous):

\[(11!) / ( (5!) (11-5)!\]

OpenStudy (anonymous):

good job!

OpenStudy (anonymous):

I mean A !

ganeshie8 (ganeshie8):

\[^{11}C_5 = \dfrac{11!}{5! (11-5)!} = \dfrac{11\times 10 \times 9 \times 8 \times 7}{5 \times 4 \times 3 \times 2 \times 1} \]

ganeshie8 (ganeshie8):

A is \(\large \color{red}{\checkmark}\)

OpenStudy (anonymous):

Thank you ! :)